# Mind the memory: Consistent time reversal removes artefactual scaling of energy dissipation rate and provides more accurate and reliable thermodynamic inference

### v3

@@ Abstract
Figure 5: (a) Tree diamond graph of depth $d=5$ with lumps of depth $l=1$ (shaded in blue) and depth $l=2$ (circled in dashed red). Transition rates in down ($+$) and up (--) direction (cyclic with respect to the thick blue edge) are taken to be $\propto\omega$ and $\propto\kappa$, respectively, chosen such that the cumulative transition rates between each level is $\omega-\kappa$. The self-similarity becomes obvious upon noticing that each tree diamond consists of multiple smaller tree diamonds. Inset: Four fundamental cycles pass through the highlighted subgraph of depth $d=2$. (b) Entropy-production estimates for different coarse-graining scales; Estimators $\dot{S}^{\rm est}_{k}$ provide an exact estimation of $\dot{S}$ for higher orders $(k\geq 2$), matching the analytical prediction (yellow line). Ignoring memory leads to a virtual energy dissipation cascade, which has no implications for dissipation on different scales. The virtual cascade is actually not even a power law, but may be mistaken for one.
Appendix C: Tree diamond graph.—We now consider a non-trivially self-similar network with a single dissipative scale (i.e., the system size; see discussion in [73]) but multiple (equivalent) dissipative cycles each with a stationary current $J_{c}$ (see Fig. 5a). It corresponds to a “tree diamond” obtained by merging two binary trees of depth $d$ such that they have the same set of leaves. Every edge is bidirectional with a rate $\propto\omega$ in $+$ and $\propto\kappa$ in -- direction, respectively, such that the cumulative transition rate between each level is $\omega-\kappa$. This way, the edge entering the root in $+$ direction has rate $\omega$, while the rates along the remaining edges are split evenly amongst the branches, (see Fig. 5a). The detailed mathematical construction is described in [73]. The exact $\dot{S}$ of the system in the cycle basis $\mathcal{C}$ with $2^{d}$ basis cycles reads (see proof in [73])
@@
We lump the network into lumps of depth ${l<d}$, whereby the $2^{d-l}$ distinct lumps on the “equator” are treated separately (see Fig. 5b). The Markovian estimate of such a lumped system is given by (see [73])
@@
Mind the memory: Consistent time reversal removes artefactual scaling of energy dissipation rate and provides more accurate and reliable thermodynamic inference
@@ I Notation
  • For $n\in\mathbb{N}$, we define $[n]\equiv\{1,\ldots n\}$.
@@ II.1.1 Filtered measure space
where $\Omega\equiv D[0,\infty)\equiv D(E\times[0,t))$ is the path space of càdlàg functions from $[0,\infty)$ to $E$. Let $\xi\equiv\mathcal{P}(E)$ be the power set of $E$, such that $(E,\xi)$ is a measurable space. The sigma-algebra $\mathcal{F}$ is on the path space defined as follows: For time $s\in[0,\infty)$ consider the evaluation mapping
@@ II.1.2 Time Reversal
where $\text{where }\theta\omega(s)$ $\equiv\omega(t-s)\text{ is the reversed path}$ (S11)
@@
and note that this exists since by Assumption 1 (i.e. thermodynamic consistency), we have absolute continuity: $\operatorname{\mathbb{P}}_{t}\ll\operatorname{\mathbb{P}}_{t}^{\theta}$ (we even equivalence. i.e. $\operatorname{\mathbb{P}}_{t}^{\theta}\ll\operatorname{\mathbb{P}}_{t}$ and hence $\operatorname{\mathbb{P}}_{t}^{\theta}\sim\operatorname{\mathbb{P}}_{t}$). By the Radon-Nikodym-Theorem, the function
@@ II.1.4 Framework
Let us denote the lumps in state space by $S_{1},\ldots,S_{o}$ ($o$ being the number of lumps). These lumps induce the subfiltration $\hat{\mathcal{F}_{t}}$ as described above. We define the index set $\mathcal{S}^{l}$ as all sequences in $[o]^{l}$ without repetition in subsequent elements
@@
~$\tilde{}\operatorname{\mathbb{P}}^{\theta}_{t,(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}(\omega)\equiv\frac{1}{\operatorname{\mathbb{P}}_{t}^{\theta}[(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}]}\operatorname{\mathbb{P}}_{t}^{\theta}(\omega),$
@@
$=$ $\sum_{(i_{1},\ldots,i_{m})}\int_{(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}\frac{{\rm d}\operatorname{\mathbb{P}}_{t}}{{\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}}(\omega){\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}\ln(\frac{1}{\operatorname{\mathbb{P}}_{t}^{\theta}[(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}]}\int_{(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}\frac{{\rm d}\operatorname{\mathbb{P}}_{t}}{{\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}}(\omega){\rm d}\operatorname{\mathbb{P}}_{t}^{\theta})$ (def. of ~$\tilde{}\operatorname{\mathbb{P}}^{\theta}_{t,\ldots}$)
@@
We note that since the underlying microscopic process is Markovian with discrete state space, we can describe the observed discrete sequence of lumps. This is exactly the embedded chain of the lumped process (which is non-Markov). When converting from a continuous time observation to the embedded chain observation—which is intrinsically discrete in time—we need to pay attention that the observed time interval was large enough. In particular, for a growing length $m$ of the observed embedded Markov chain, the continuous in time observation interval $[0,t)$ needs to grow sufficiently fast for enough jumps to be observed. It cannot grow too fast either, since otherwise we discard too many discrete jumps, and the lower bound we would get would converge to 00.
@@
×$\times\operatorname{\mathbb{P}}\left[{\hat{n}(t)\geq m(t)}\right]$
@@
×$\times\operatorname{\mathbb{P}}\left[{\hat{n}(t)\geq m(t)}\right].$ (by assumption, the lumped process has order $k$)
@@
$\operatorname{\mathbb{E}}\left[N(t)\right]=$ $\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]missing}+\mathcal{O}(1)$
@@
$\leq$ [|$\operatorname{\mathbb{P}}\left[\absolutevalue{N(t)-\operatorname{\mathbb{E}}\left[N(t)\right]missing}>t^{2/3}\right]$
@@
where we used Chebyshev’s inequality in the last step. Here, $m(t)=\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]missing}-o\left(t\right)$. If we instead consider a chain consisting of only the fastest lumps, we would have $m(t)=\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{fast}}\right]missing}-o\left(t\right)$. Thus, there is a constant $\tau$ such that $\operatorname{\mathbb{E}}\left[\tau_{\text{fast}}\right]<\tau<\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]$, so that with $m(t)=\frac{t}{\tau}-o\left(t\right)$ we have convergence. For the particular choice of $\tau$, note that $\hat{n}(t)$ will, by the central limit theorem (recall that we have a narrow distribution), concentrate around its mean. Therefore, the constant $\tau$ will be the average time per lump.
@@
so that our $k$-th order estimator provides—given that the order of the process is $k$—always a guaranteed lower bound. ∎
@@ II.2 Robustness of the estimator
$={\operatorname{\mathbb{P}}\left[{X^{(k+2)}=\hat{\gamma}_{k}^{(k+1)}\mid X^{(2:k+1)}\hat{\gamma}_{k}^{(1:k)}}\right]},$
@@
$=\frac{1}{T}\sum_{\hat{\gamma}_{k}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(k+1)}\mid\hat{\gamma}_{k}^{(k-n+1:k)}}\right]}}-\frac{1}{T}\sum_{\hat{\gamma}_{k}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1)}\mid\hat{\gamma}_{k}^{(n+1:2)}}\right]}}$ )
$=\frac{1}{T}\sum_{\hat{\gamma}_{k}^{(k-n+1:k+1)}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(k-n+1:k+1)}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(k+1)}\mid\hat{\gamma}_{k}^{(k-n+1:k)}}\right]}}$ )
$-\frac{1}{T}\sum_{\hat{\gamma}_{k}^{(1:n+1)}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1:n+1)}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1)}\mid\hat{\gamma}_{k}^{(n+1:2)}}\right]}}.$ ) . (marginalisation)
@@
$=\frac{1}{T}\sum_{\hat{\gamma}_{k}^{(1:n+1)}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1:n+1)}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(n+1)}\mid\hat{\gamma}_{k}^{(1:n)}}\right]}}-\frac{1}{T}\sum_{\hat{\gamma}_{k}^{(1:n+1)}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1:n+1)}}\right]}\ln{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1)}\mid\hat{\gamma}_{k}^{(n+1:2)}}\right]}}$ )
$=\frac{1}{T}\sum_{\hat{\gamma}_{k}^{(1:n+1)}}{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1:n+1)}}\right]}\ln{\frac{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(n+1)}\mid\hat{\gamma}_{k}^{(1:n)}}\right]}}{{\operatorname{\mathbb{P}}\left[{\hat{\gamma}_{k}^{(1)}\mid\hat{\gamma}_{k}^{(n+1:2)}}\right]}}}$ )

### v4

@@ Abstract
Figure 6: (a) Tree diamond graph of depth $d=5$ with lumps of depth $l=1$ (shaded in blue) and depth $l=2$ (circled in dashed red). Transition rates in down (++) and up (-) direction (cyclic with respect to the thick blue edge) are taken to be $\propto\omega$ and $\propto\kappa$, respectively, chosen such that the cumulative transition rates between each level is $\omega-\kappa$. The self-similarity becomes obvious upon noticing that each tree diamond consists of multiple smaller tree diamonds. Inset: Four fundamental cycles pass through the highlighted subgraph of depth $d=2$. (b) Entropy-production estimates for different coarse-graining scales; Estimators $\dot{S}^{\rm est}_{k}$ provide an exact estimation of $\dot{S}$ for higher orders $(k\geq 2$), matching the analytical prediction (yellow line). Ignoring memory leads to a virtual energy dissipation cascade, which has no implications for dissipation on different scales. The virtual cascade is actually not even a power law, but may be mistaken for one.
Appendix C: Tree diamond graph.—We now consider a non-trivially self-similar network with a single dissipative scale, the system size (see discussion in [72]), but multiple (equivalent) dissipative cycles each with a stationary current $J_{c}$ (see Fig. 6a). It corresponds to a “tree diamond” obtained by merging two binary trees of depth $d$ such that they have the same set of leaves. Every edge is bidirectional with a rate $\propto\omega$ in ++ and $\propto\kappa$ in - direction, respectively, such that the cumulative transition rate between each level is $\omega-\kappa$. This way, the edge entering the root in ++ direction has rate $\omega$, while the rates along the remaining edges are split evenly amongst the branches, (see Fig. 6a). The detailed mathematical construction is described in [72]. The exact $\dot{S}$ of the system in the cycle basis $\mathcal{C}$ with $2^{d}$ basis cycles reads (proof in [72])
@@
We lump the network into lumps of depth l<d{l<d}, whereby the $2^{d-l}$ distinct lumps on the “equator” are treated separately (see Fig. 6b). The Markovian estimate of such a lumped system is given by (see [72])
@@
Appendix D: Nonexistence of scaling of $\dot{S}_{\rm M}$ under time-scale separation.—We already showed that a power-law scaling of $\dot{S}_{\rm M}$ has no physical implications in absence of a time-scale separation due to inconsistent time reversal. In the presence of a time-scale separation the time reversal becomes consistent [29], but the power-law scaling disappears, as we now show.
Suppose there is a time-scale separation, i.e. suppose the dynamics within each element (lump) of a lumping $U$ relaxes fast enough for the transitions between elements of $U$ to proceed as in Ref. [29] and (consistently) apply a Markovian estimator $\dot{S}^{\rm est}_{1}$. We now argue that time-scale separation in $U$ implies time-scale separation in each refined scale. Consider $W$ to be a refinement of $U$  [^2] That is, $W$ is a lumping that refines the partition $U$.. Due to time-scale-separation in $U$ any transition between and within lumps of $W$ happens at least as fast as in $U$. Hence, we also have time-scale separation in $W$.
Hence, in presence of time-scale separation, the Markovian time reversal is consistent, and the Markovian estimate gives the exact result [29] for scale $U$ and any refined scale. This naturally implies the absence of (power law) scaling of $\dot{S}_{1}^{\mathrm{e}st}$.
Figure 7: (a) Cycle graph $C_{36}$ with transition rates chosen as indicated in the inset and detailed in [72] (Section XI.) such that the dynamics exhibit time-scale separation. (b) Entropy-production rate estimates for $C_{36}$ with time-scale separation between lumps of size 66: The time-scale separation assumption implies that any refined scale is Markovian ($\dot{S}_{1}^{est}=\dot{S}_{2}^{est}$). This also implies, however, that there is no scaling of the entropy production rate (observation).
To illustrate this, consider the cycle graph $C_{36}$ with rates chosen to have time-scale separation (the detailed parameters are given in Fig. S7 in [72]). At a lumping of size $6$ (first lump are vertices $1,\ldots,6$, second lump vertices $7,\ldots,12$ etc.), we observe a time-scale separation. Thus, we may use a Markovian estimator for a lumping of size $6$. To get the correct entropy production rate at a lumping of size $12$, however, we must use an estimator $\dot{S}_{k}^{\rm est}$ of order $k\geq 2$ (as argued in the letter). As we see in Fig. 7, any refinement of the lumping of size $6$ will be Markovian and result in the same estimate. Thus, there is no power-law scaling under time-scale separation.
Appendix E: Uncountability of microscopic paths.—We finally show that the number of microscopic paths projecting onto a mesoscopic path in any time window $[0,t]$ (with $t>0$) is generally not countable. To that end, suppose that a mesostate $M$ contains at least two neighboring microscopic states $\{0,1\}$. Take some real number $r\in[0,1]$, and consider its binary encoding, say $\operatorname{bin}(r)$. Take the corresponding microscopic trajectory of jumps between states $0,1$. Furthermore, a corresponding sequence of increasing jumptimes in $[0,t]$ exists. This is a unique mapping to a microscopic path, but all those paths correspond to the mesoscopic path of staying in $M$. By uncountability of $\mathbb{R}$, the claim follows.
Appendix F: Other parameter choices.— We present here how other choices of $u,v$ in Fig. 1 affect the estimators’ quality: Any $k\geq 2$ is better than the Markovian estimate $\dot{S}_{\rm M}=\dot{S}^{\mathrm{est}}_{1}$ [49]. We always observe saturation for $k\geq 3$ (upto statistical fluctuations) and our estimate is already at $k=2$ better than $\frac{1}{2}\dot{S}^{\rm uncond}_{2}/\dot{S}$ [51]. Note how saturation happens for $u=1$ already at $k\geq 2$, indicating that at $u=1$ we have an order-2 process: While $\dot{S}^{\mathrm{est}}_{2}$ allows for $100\%$ recovery, using in that setting $\frac{1}{2}\dot{S}^{\rm uncond}_{2}/\dot{S}$ [51] would allow only for recovery of $82\%$ and $83\%$ respectively, with the remaining recovery only possible by through the statistically very costly (binning of a continuous-time random variable for integral computation) inference from the waiting time distribution asymmetry. Note that $\sigma_{2}^{\ell}$ in Ref. [51] is—in contrast to $\sigma_{2}^{t}$—statistically equally feasible as $\dot{S}_{2}^{\rm est}$.
parameter u 2.0 1.0 2.0 1.0 1.5
parameter v 2.0 2.0 1.0 3.0 2.0
$\dot{S}^{\mathrm{est}}_{1}/\dot{S}$ [49] 0.03 0.58 0.06 0.63 0.14
$\dot{S}^{\mathrm{est}}_{2}/\dot{S}$ 0.21 1.05 0.09 1.03 0.45
$\frac{1}{2}\dot{S}^{\rm uncond}_{2}/\dot{S}$ [51] 0.12 0.82 0.07 0.83 0.30
$\dot{S}^{\mathrm{est}}_{3}/\dot{S}$ 0.39 1.02 0.20 1.02 0.61
$\dot{S}^{\mathrm{est}}_{4}/\dot{S}$ 0.41 1.02 0.22 1.02 0.63
$\dot{S}^{\mathrm{est}}_{5}/\dot{S}$ 0.41 1.00 0.21 1.01 0.62
$\dot{S}^{\mathrm{est}}_{6}/\dot{S}$ 0.4 0.99 0.21 1.00 0.61
Table 1: Ability of different estimators under more parameter choices for the system of Fig. 1: We observe a consistently better approximation of $\dot{S}$ through $\dot{S}_{k}^{\rm est}$ for every $k\geq 2$ compared to $\dot{S}_{M}=\dot{S}_{1}^{\rm est}$ [49] and compared to $\dot{S}_{2}^{\rm uncond}/2=\frac{1}{2}(\dot{S}_{1}^{\rm est}+\dot{S}_{2}^{\rm est})=\sigma_{2}^{\ell}$ in [51].
@@
Consistent time reversal and reliable and accurate inference in the presence of memory
@@ I Notation
  • For $n\in\mathbb{N}$, we define $n]\equiv\{1,\ldots n\}$.
@@ II.1.1 Filtered measure space
where $\Omega\equiv D[0,\infty)\equiv D(E\times[0,t))$ is the path space of càdlàg functions from $0,\infty)$ to $E$. Let $\xi\equiv\mathcal{P}(E)$ be the power set of $E$, such that $(E,\xi)$ is a measurable space. The sigma-algebra $\mathcal{F}$ is on the path space defined as follows: For time $s\in[0,\infty)$ consider the evaluation mapping
@@ II.1.2 Time Reversal
$\text{where }\theta\omega(s)$ $\equiv\omega(t-s)\text{ is the reversed path}$ (S11)
@@
and note that this exists since by Assumption 1 (i.e. thermodynamic consistency), we have absolute continuity: $\operatorname{\mathbb{P}}_{t}\ll\operatorname{\mathbb{P}}_{t}^{\theta}$ (we even have equivalence, i.e. $\operatorname{\mathbb{P}}_{t}^{\theta}\ll\operatorname{\mathbb{P}}_{t}$ and hence $\operatorname{\mathbb{P}}_{t}^{\theta}\sim\operatorname{\mathbb{P}}_{t}$). By the Radon-Nikodym-Theorem, the function
@@ II.1.4 Framework
Let us denote the lumps in state space by $S_{1},\ldots,S_{o}$ ($o$ being the number of lumps). These lumps induce the subfiltration $\hat{\mathcal{F}_{t}}$ as described above. We define the index set $\mathcal{S}^{l}$ as all sequences in $o]^{l}$ without repetition in subsequent elements
@@
One might then attempt to apply the log-sum inequality to the embedded Markov chain. However, as discussed earlier, evaluating the probability that all microscopic trajectories $\gamma$ corresponds to a specific mesoscopic trajectory $\hat{\gamma}_{k}$ requires dealing with an uncountably infinite series of integrals over the time interval $[0,t]$ (see Appendix E in the main Letter). Hence, we require the measure-theoretic approach introduced above.
@@
$\tilde{}\operatorname{\mathbb{P}}^{\theta}_{t,(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}(\omega)\equiv\frac{1}{\operatorname{\mathbb{P}}_{t}^{\theta}[(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}]}\operatorname{\mathbb{P}}_{t}^{\theta}(\omega),$
@@
We are now in a position to employ the measure theoretic version of Jensen’s inequality in path space. This allows us to cope with the challenges of continuous time.
@@
$=$ $\sum_{(i_{1},\ldots,i_{m})}\int_{(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}\frac{{\rm d}\operatorname{\mathbb{P}}_{t}}{{\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}}(\omega){\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}\ln(\frac{1}{\operatorname{\mathbb{P}}_{t}^{\theta}[(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}]}\int_{(S_{i_{1}},\ldots,S_{i_{m}})_{t}^{*}}\frac{{\rm d}\operatorname{\mathbb{P}}_{t}}{{\rm d}\operatorname{\mathbb{P}}_{t}^{\theta}}(\omega){\rm d}\operatorname{\mathbb{P}}_{t}^{\theta})$ (def. of $\tilde{}\operatorname{\mathbb{P}}^{\theta}_{t,\ldots}$)
@@
We note that since the underlying microscopic process is Markovian with discrete state space, we can describe the observed discrete sequence of lumps. This is exactly the embedded chain of the lumped process (which is non-Markov). When converting from a continuous time observation to the embedded chain observation—which is intrinsically discrete in time—we need to pay attention that the observed time interval was large enough. In particular, for a growing length $m$ of the observed embedded Markov chain, the continuous in time observation interval $[0,t)$ needs to grow sufficiently fast for enough jumps to be observed. It cannot grow too fast either, since otherwise we discard too many discrete jumps, and the lower bound we would get would converge to 0.
@@
$\times\operatorname{\mathbb{P}}\left[{\hat{n}(t)\geq m(t)}\right]$
@@
$\times\operatorname{\mathbb{P}}\left[{\hat{n}(t)\geq m(t)}\right].$ (by assumption, the lumped process has order $k$)
@@
$\operatorname{\mathbb{E}}\left[N(t)\right]=$ $\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]}+\mathcal{O}(1)$
@@
[…]
@@
where we used Chebyshev’s inequality in the last step. Here, $m(t)=\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]}-o\left(t\right)$. If we instead consider a chain consisting of only the fastest lumps, we would have $m(t)=\frac{t}{\operatorname{\mathbb{E}}\left[\tau_{\text{fast}}\right]}-o\left(t\right)$. Thus, there is a constant $\tau$ such that $\operatorname{\mathbb{E}}\left[\tau_{\text{fast}}\right]<\tau<\operatorname{\mathbb{E}}\left[\tau_{\text{slow}}\right]$, so that with $m(t)=\frac{t}{\tau}-o\left(t\right)$ we have convergence. For the particular choice of $\tau$, note that $\hat{n}(t)$ will, by the central limit theorem (recall that we have a narrow distribution), concentrate around its mean. Therefore, the constant $\tau$ will be the average time per lump.
@@
so that our $k$-th order estimator provides—given that the order of the process is kk—always a guaranteed lower bound. ∎
@@ II.2 Robustness of the estimator
$={\operatorname{\mathbb{P}}\left[{X^{(k+2)}=\hat{\gamma}_{k}^{(k+1)}\mid X^{(2:k+1)}=\hat{\gamma}_{k}^{(1:k)}}\right]},$
@@
[…]
@@
[…]